OpenStudy (anonymous):

One solution contains 50% alcohol and another solution contains 80% alcohol. How many liters of each solution should be mixed to produce 10.5 liters of a 70% alcohol situation?

OpenStudy (amistre64):

x(.5) + y(.8) = 10.5(.7) x+y=10.5 x=(10.5-y) (10.5 -y)(.5) + .8y = 7.35 5.25 -.5y +.8y = 7.35 .3y = 7.35 - 5.25 .3y = 2.1 y= (2.1)/(.3) = 21/3 = 7 x+7 = 10.5 x= 3.5 Double check: 3.5(.5) + 7(.8) = 10.5(.7) 1.75 + 5.6 = 7.35 7.35 = 7.35 (GOOD) x= 3.5(50%); and y=7(80%)